# EVQAScore: Efficient Video Question Answering Data Evaluation

### v2

@@ Abstract
Video question-answering (QA) is a core task in video understanding. Evaluating the quality of video QA and video caption data quality for training video large language models (VideoLLMs) is an essential challenge. Although various methods have been proposed for assessing video caption quality, there remains a lack of dedicated evaluation methods for Video QA. To address this gap, we introduce EVQAScore, a reference-free method that leverages keyword extraction to assess both video caption and video QA data quality. Additionally, we incorporate frame sampling and rescaling techniques to enhance the efficiency and robustness of our evaluation, this enables our score to evaluate the quality of extremely long videos. Our approach achieves state-of-the-art (SOTA) performance (32.8 for Kendall correlation and 42.3 for Spearman correlation, 4.7 and 5.9 higher than the previous method PAC-S++) on the VATEX-EVAL benchmark for video caption evaluation. Furthermore, by using EVQAScore for data selection, we achieved SOTA results with only 12.5% of the original data volume, outperforming the previous SOTA method PAC-S and 100% of data. Our code is available at https://anonymous.4open.science/r/EVQAScore-FFA6/.
@@ 1 Introduction
With the rapid advancements in large language models (LLMs) [30, 36] and multimodal large language models (MLLMs) [45, 40],
Figure 1: We give an example of EVQAScore, from the figure we can see that Answer1 is low quality and contains few information while Answer2 contains more information and is a more comprehensive answer. We can see our EVQAScore is the only score can correctly identify the high quality data.
@@
Among MLLMs, VideoLLMs achieve competitive performance in traditional multimodal tasks such as visual recognition [44], video understanding [41, 35], and action recognition [39]. Moreover, their excellent language understanding capabilities enable strong performance in text-rich tasks, such as video question-answering [9] and video-centric dialogues [39].
@@
Most existing VideoLLMs focus on modifying model architectures to effectively utilize information from multiple modalities [38, 39, 26, 43, 17]. While model effectiveness is crucial, data quality also plays a significant role in the success of VideoLLMs. For instance, Li et al. [18], Wang et al. [39] demonstrate that higher-quality training data enhances the performance of VideoLLMs. Therefore, evaluating video QA data quality is essential. However, previous methods such as EMScore and PAC-S [34, 33] focus exclusively on video caption evaluation while neglecting video QA evaluation. Moreover, these methods struggle with understanding the correct semantics of longer captions or QA pairs. As illustrated in Figure 1, EMScore and PAC-S fail to identify higher-quality video QA pairs. These methods overlook the importance of efficient and robust data management and face three key challenges:
C1. Inability to Accurately Understand Caption and QA Semantics Previous methods such as EMScore [34] and PAC-S [33] rely on TF-IDF to evaluate video caption data. However, TF-IDF merely computes word frequency and does not capture semantic meaning. As a result, these methods often fail to accurately evaluate video captions and are unable to assess video QA data with precision. Thus, there is a need for improved semantic understanding in Video QA evaluation.
C2. Low Evaluation Efficiency Given the extensive volume and length of video data, large-scale training datasets frequently occupy substantial storage space, ranging from several hundred gigabytes to tens of terabytes [38, 39, 20]. Evaluating video caption and QA data is time-consuming and computationally costly. Previous methods like EMScore [34] and PAC-S [33] process every frame, leading to considerable computational waste. Therefore, developing a computationally efficient evaluation metric is essential.
C3. Lack of VideoLLM Experiments Data plays a crucial role in the success of VideoLLMs [38, 39, 26, 43, 17]. Although various methods have been proposed to evaluate the quality of video caption data, they have not incorporated VideoLLM experiments to assess the effectiveness of their methods in training VideoLLMs. Consequently, additional VideoLLM experiments are needed.
@@
To address these challenges, we introduce EVQAScore, which utilizes LLMs to extract keywords for enhanced video captioning and video QA data evaluation. To support the evaluation of longer videos, we efficiently compute EVQAScore through uniform frame sampling, which reduces computational costs by a factor of 30 without affecting results. Compared to previous methods such as EMScore [34] and PAC-S [33], our approach can be applied to video QA evaluation and achieves superior visual understanding through keyword extraction, outperforming the TF-IDF technique used by EMScore and PAC-S. Additionally, due to its efficiency and keyword extraction capabilities, our method can handle extremely long captions and videos, a feat not possible for EMScore and PAC-S because of the 77-word limit imposed by CLIP. The core contributions of this paper are summarized as follows:
@@
  • New Method: We introduce a novel approach, EVQAScore, designed for the evaluation of Video QA and video caption data quality. To the best of our knowledge, this is the first systematic effort to evaluate the quality of Video QA data. Our approach leverages LLMs and keyword extraction techniques to improve the understanding of QA data in Video QA evaluation. Additionally, we employ uniform frame sampling to enhance efficiency and scalability, allowing EVQAScore to be applied to extremely long videos and large datasets.
  • Efficient Video QA Data Evaluation: We use uniform frame extraction, resulting in a computational cost reduction of over 30 times compared to processing all frames. Furthermore, we demonstrate that this reduction does not lead to any performance degradation.
@@
  • SOTA Performance for Video Caption Evaluation: We utilize LLM for keyword extraction for better caption and QA understanding. We evaluated the performance of our method on the VATEX-Eval benchmark for Video Caption data, achieving a Kendall correlation of 32.8 and a Spearman correlation of 42.3, which are 4.7 and 5.9 points higher than the previous method PAC-S++.
@@
  • SOTA Performance for Video QA Evaluation: We constructed a dataset comprising 400K video entries, consisting of 200K noisy samples and 200K high-quality samples. Using our EVQAScore and the previous SOTA model PAC-S, we filtered the data, retrieving a significantly lower amount of noisy data than PAC-S. As shown in Figure 1, our method is the only method that can identify low-quality video QA data. Additionally, when evaluated on ActivityNet, MSRVTT, MSVD, TGIF, MVBench, and two longer benchmarks—VideoChatGPT Bench and VideoChatGPT Bench Diverse—our EVQAScore outperformed PAC-S and 100% of data with only 12.5% of data.
@@ 2 Related Work
2.1 Video Data Evaluation
Traditional image and video captioning evaluations have relied on metrics like BLEU[32], METEOR[4], ROUGE[21], CIDEr[37], and SPICE[1]. These metrics, based on n-gram matching, primarily focus on lexical similarity between generated and reference captions. While effective in text-based tasks like machine translation, they struggle to capture the semantic richness of video content, leading to poor correlation with human judgments, especially when evaluating dynamic, time-sensitive visual elements in videos.
@@ 2.1 Video Data Evaluation
2.2 Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data [36, 31]. In this scenario, the quality and quantity of data become paramount [3]. In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models [29]. As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation [3]. Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection [3]. For instance, Du et al. [6] leverages DeBERTa [7] for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. [5] score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. [42] use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality.
@@ 2.2 Data Quality and Selection
Recently, inspired by the remarkable understanding capabilities of LLMs [14], multimodality breakthroughs [14, 15, 11, 12, 16, 13, 25, 22, 10], and pre-trained models, researchers have started using LLMs to understand videos, achieving SoTA results [18, 43, 2, 39, 20]. VideoLLaMA [43] is one of the pioneering studies in VideoLLMs, utilizing a visual encoder and a video Q-Former projector to understand videos. However, due to its Q-Former structure, the computational cost is high. To address this, subsequent works [20] adopted the LLaVA [24, 23] MLP structure, significantly reducing computational costs while still achieving SoTA performance. Similarly, MiniGPT4Video [2] uses an MLP adapter for efficient training.
@@ 3 Method
Figure 2: Illustration of the computation of EMScore which utilize coarse to fine video evaluation. We utilize keyword extraction and then use frames and keywords as ground truth for precision and recall calculation.
3.1 Motivation and problem formulation
Inspired by EMScore and PAC-S, we extend these video-caption data quality evaluation methods to assess the quality of VideoQA data. We aim to address three key limitations of existing methods.
Difficulty in Evaluating Long Videos:
Previous methods, such as PAC-S, face challenges when dealing with longer videos due to the high computational costs and unable to understand long captions and QAs due to the CLIP token limitation (77 tokens). To address this issue, we employ uniform frame extraction to reduce computational costs and utilize keyword extraction for better QA and caption understanding.
Inability to Capture Accurate Semantic Meaning of Captions:
Methods like PAC-Score and EMScore rely on TF-IDF for semantic understanding, which is limited because it only computes word frequency and cannot capture contextual semantics. Therefore, we utilize LLMs to extract keywords for more accurate understanding of QA and caption semantics.
Evaluating VideoQA Data for VideoLLMs:
Video QA is increasingly important in areas such as SFT, requiring adaptable methods capable of handling these tasks. As such, we apply EVQAScore for Video QA data filtering.
Problem Formulation
Given a video $V$ and a corresponding question-answer pair $(Q,A)$, our goal is to evaluate the quality of the VideoQA data using our proposed evaluation metric, EVQAScore. This metric assesses the data’s suitability for training VideoLLMs by scoring based on several criteria, including semantic relevance and informativeness of the answer $A$ with respect to the video $V$ and the question $Q$.
Formally, for a given data instance $(V,Q,A)$, we compute EVQAScore to evaluate the quality of $(V,Q,A)$:
@@ 3.1 Motivation and problem formulation
Our objective is to filter and select high-quality data based on the EVQAScore to improve the performance and training efficiency of VideoLLMs.
3.2 Efficiency of EVQAScore
Keyframe Extraction
To address the challenge of evaluating long videos efficiently, we adopt a uniform frame sampling strategy to reduce the number of frames processed while retaining critical information. By sampling keyframes at regular intervals, we capture essential content and minimize computational costs. The frame sampling process is defined as:
$S=\{v_{i}\mid i=k\cdot l,\;k\in\mathbb{Z}_{\geq 0},\;k\cdot l\leq m\}$ (1)
where $S$ represents the set of sampled frames, $v_{i}$ denotes the $i$-th frame in the video $V$, $m$ is the total number of frames, and $l$ is the sampling interval. This approach reduces redundancy while ensuring a diverse representation of keyframes.
In our experiment we select sampling interval $l$ to be 10, 20, and 30. We present the result in Table 1.
@@ 3.2 Efficiency of EVQAScore
Given a text input $X$, we use a large language model (LLM) to extract a list of key phrases $\{k_{1},k_{2},\ldots,k_{n}\}$ as follows:
@@
$K=\{k_{1}\oplus k_{2}\ldots\oplus k_{n}\}$
where $\text{LLM}_{\text{extract}}$ denotes the language model’s extraction function, and $P$ represents the prompt designed to guide the extraction of relevant phrases from the input $X$. This approach enhances the efficiency of video-text alignment by focusing on the most relevant key phrases identified through the prompt-based extraction.
3.3 Video caption evaluation
The improved scoring method now combines the coarse-grained score and the modified fine-grained score, which includes keyframe-based matching and keyword-based matching. This enhanced PAC-Score improves efficiency when handling long videos and complex text descriptions, focusing on the most essential frames and phrases.
@@ 3.3 Video caption evaluation
$P(V,X)_{f}=P\left(S,K\right)_{f}=\frac{1}{|X|}\sum_{k_{j}\in K}\max_{v_{i}\in S}f_{v_{i}}^{\top}f_{k_{j}}$
$R(V,X)_{f}=R\left(S,K\right)_{f}=\frac{1}{|V|}\sum_{v_{i}\in S}\max_{k_{j}\in K}f_{v_{i}}^{\top}f_{k_{j}}$
where $f_{v_{i}}$ and $f_{k_{j}}$ represent the CLIP embedding scores of frame $v_{i}$ and keyword $k_{j}$. Subsequently, the fine-grained EVQAScore is calculated as:
@@
Finally, the overall EVQAScore is computed by combining the original coarse-grained score with the updated fine-grained score, which incorporates both keyframe and keyword matching.
@@
This formula ensures that both visual and textual representations are aligned more effectively, improving precision, recall, and computational efficiency by focusing on keyframes and keywords.
3.4 Video QA Evaluation
We compute EVQAScore for Video QA data using the following formula:
$\text{EVQAScore}(\text{Concat}(Q,A),V)\cdot\log(n+1)$
In this formulation, $\text{Concat}(Q,A)$ represents the concatenation of the question $Q$ and the answer $A$, while $n$ denotes the number of extracted keywords. We posit that a higher number of extracted keywords often indicates a richer and more informative caption, thereby warranting a higher score. We show the advancement of our EVQAScore in Section 4.
Table 1: Comparison of time spent managing video data between uniform keyframe extraction (every 10 frames) and no keyframe extraction for 1000 videos. Extracting keyframes leads to a significantly shorter time.

### v3

# EVQAScore: A Fine-grained Metric for Video Question Answering Data Quality Evaluation

@@ Abstract
Video question-answering (QA) and video-caption data play a central role in video understanding. Evaluating the quality of both VideoQA and video-caption data for training video large language models (VideoLLMs) presents a significant challenge. While several methods have been proposed for assessing video-caption quality, there is a lack of dedicated evaluation methods specifically for VideoQA. To address this gap, we introduce EVQAScore, a reference-free method that employs patch retrieval for fine-grained evaluation of video-caption and VideoQA data quality. Additionally, we utilize a keyword extraction method to evaluate longer captions (previous methods like PAC-S and EMScore are limited by CLIP’s 77-word constraint) and to improve semantic understanding. Furthermore, we incorporate frame sampling to enable faster evaluation without sacrificing performance. Our approach achieves state-of-the-art (SOTA) performance on the VATEX-EVAL benchmark for video-caption evaluation, with scores of 28.9 for Kendall correlation and 37.4 for Spearman correlation, surpassing the previous method PAC-S++ by 0.8 and 1.0 points, respectively. By using EVQAScore for data selection, we achieve SOTA results with only 12.5% of the original data volume, outperforming PAC-S, which required 100% of the data. Our code is available at https://anonymous.4open.science/r/EVQAScore-FFA6/.
@@ 1 Introduction
Figure 1: We provide an example of the EVQAScore evaluation process. In (a), we sample frames and apply YOLO for patch recognition. In (b), we extract keywords from the question-answer pairs. Finally, we calculate the EVQAScore based on the precision and recall between patches and keywords.
@@
With the rapid advancements in large language models (LLMs) [25, 31] and multimodal large language models (MLLMs) [39, 35], VideoLLMs have achieve competitive performance in traditional multimodal tasks such as visual recognition [38], video understanding [36, 30], and action recognition [34]. Moreover, their excellent language understanding capabilities enable strong performance in text-rich tasks, such as video question-answering [5] and video-centric dialogues [34].
@@
Most existing VideoLLMs focus on modifying model architectures to effectively leverage information from multiple modalities [33, 34, 22, 37, 13]. While model effectiveness is critical, data quality also plays a pivotal role in the success of VideoLLMs. For example, Li et al. [14], Wang et al. [34] demonstrate that higher-quality training data significantly improves the performance of VideoLLMs. Therefore, evaluating the quality of VideoQA data is essential. On the other hand, previous methods, such as EMScore and PAC-S [29, 28], have focused on evaluating human-preference video-captions. In this work, we aim to develop an evaluation metric that not only excels in human-preference evaluation but also provides high-quality training data for VideoLLMs. We set the following three tasks as our goals for improvement.
Detailed Information from Video Frames
Existing methods (PAC-S and EMScore) typically use a single frame as the minimum unit for video feature extraction, potentially overlooking the rich information within a single frame. To address this limitation, we employ YOLO [27] to extract key objects from video frames, enabling the utilization of diverse information within a single frame.
Better Semantic Understanding
Methods such as PAC-Score [28] and EMScore [29] rely on TF-IDF for semantic understanding, which is limited as it only computes word frequency and fails to capture contextual semantics. To address this limitation, we employ LLMs to extract keywords, enabling a more accurate understanding of both QA and caption semantics. Furthermore, in the context of long videos with extended captions and QAs, which often exceed 77 words (with over 30% of the VCG+ [24] Train Dataset containing examples longer than 77 words), PAC-Score and EMScore become less effective. As a result, there is a need for an improved evaluation method capable of supporting longer videos and arbitrarily lengthy captions and QAs, providing better semantic comprehension.
VideoQA Data Quality Evaluation
As VideoQA becomes increasingly important in areas such as SFT, evaluating the quality of VideoQA data is crucial. Therefore, we need a evaluation method supporting both video-caption and VideoQA quality evaluation.
@@
To improvement the from the three aspects, we introduce EVQAScore, which utilizes LLMs to extract keywords for enhanced caption and QA data understanding. Moreover, we use YOLO [27] to extract fine-grained feature. To support the evaluation of longer videos, we need to control the computational cost. Therefore we efficiently compute EVQAScore through uniform frame sampling, which reduces computational costs by a factor of 30 without affecting results. Compared to previous methods such as EMScore [29] and PAC-S [28], our approach can be applied to videoQA evaluation and achieves superior visual understanding through keyword extraction, outperforming the TF-IDF technique used by EMScore and PAC-S. Additionally, due to its efficiency and keyword extraction capabilities, our method can handle extremely long captions and QAs for videos, a feat not possible for EMScore and PAC-S because of the 77-word limit imposed by CLIP.
The core contributions of this paper are summarized as follows:
@@
  • New Method: We introduce a novel approach, EVQAScore, designed for the evaluation of VideoQA and video-caption data quality. To the best of our knowledge, this is the first systematic effort to evaluate the quality of VideoQA data. Our approach leverages LLMs and keyword extraction techniques to improve the understanding of QA data in VideoQA evaluation. Additionally, we employ uniform frame sampling to enhance efficiency and scalability, allowing EVQAScore to be applied to extremely long videos and large datasets.
@@
  • SOTA Performance for Video-Caption Evaluation: We utilize LLM for keyword extraction for better caption and QA understanding. We evaluated the performance of our method on the VATEX-Eval benchmark for video-caption data, achieving a Kendall correlation of 28.9 and a Spearman correlation of 37.4, which are 0.8 and 1.0 points higher than the previous method PAC-S++.
@@
  • SOTA Performance for Filtering VideoLLMs Training Data: We constructed a dataset comprising 400K video entries, consisting of 200K noisy samples and 200K high-quality samples. Using our EVQAScore and the previous SOTA model PAC-S, we filtered the data, retrieving a significantly lower amount of noisy data than PAC-S. As shown in Figure 1, our method is the only method that can identify low-quality VideoQA data. Additionally, when evaluated on ActivityNet, MSRVTT, MSVD, TGIF, MVBench, and two longer benchmarks—VideoChatGPT Bench and VideoChatGPT Bench Diverse—our EVQAScore outperformed PAC-S and 100% of data with only 12.5% of data.
@@ 2 Related Work
2.1 Video-Caption Data Quality Evaluation
Video Data Quality Evaluation
Traditional image and video-captioning evaluations have relied on metrics like BLEU[26], METEOR[3], ROUGE[17], CIDEr[32], and SPICE[1]. These metrics, based on n-gram matching, primarily focus on lexical similarity between generated and reference captions. While effective in text-based tasks like machine translation, they struggle to capture the semantic richness of video content, leading to poor correlation with human judgments, especially when evaluating dynamic, time-sensitive visual elements in videos.
Video-Caption Data Quality Evaluation
@@ 2.1 Video-Caption Data Quality Evaluation
Recently, inspired by the remarkable understanding capabilities of LLMs [10], multimodality breakthroughs [10, 11, 7, 8, 12, 9, 21, 18, 6], and pre-trained models, researchers have started using LLMs to understand videos, achieving SoTA results [14, 37, 2, 34, 16]. VideoLLaMA [37] is one of the pioneering studies in VideoLLMs, utilizing a visual encoder and a video Q-Former projector to understand videos. However, due to its Q-Former structure, the computational cost is high. To address this, subsequent works such as VideoLLaVA [16], LLaVA-Next [20] adopted the LLaVA [20, 19] MLP structure, significantly reducing computational costs while still achieving SoTA performance. Similarly, MiniGPT4Video [2] uses an MLP adapter for efficient training.
@@ 3 Method
Figure 2: We illustrate the computation of $\text{EVQAScore}_{f}$. First, fine-grained patches are extracted from the sampled frames. Next, keyword extraction is performed to enhance caption and question-answer semantic understanding. Finally, the patches and keywords are combined to calculate precision and recall, resulting in the EVQAScore.
3.1 Problem Formulation
Inspired by EMScore and PAC-S, we extend these video-caption data quality evaluation methods to assess both video-caption and VideoQA data quality with a more fine-grained metric, EVQAScore. We now present the problem formulation:
Given a video $V$ and a corresponding question-answer pair $(Q,A)$, our goal is to evaluate the quality of the VideoQA data using the proposed EVQAScore metric. This metric evaluates the data’s suitability for training VideoLLMs by scoring based on several criteria, including the semantic relevance and informativeness of the answer $A$ in relation to the video $V$ and the question $Q$.
Formally, for a given data instance $(V,Q,A)$, where $V=\{v_{i}\}_{i=1}^{|V|}$ ($|V|$ is the number of video frames), we compute the EVQAScore to assess the quality of $(V,Q,A)$:
@@ 3.1 Problem Formulation
Our objective is to filter and select high-quality data based on the EVQAScore to 1) select data aligned with human preferences, and 2) improve the performance and training efficiency of VideoLLMs.
3.2 Data Preparation for EVQAScore
To address the challenge of evaluating long videos efficiently, we adopt a uniform frame sampling strategy to reduce the number of frames processed while retaining critical information.
Frame Sampling
By sampling frames at regular intervals, we capture essential content and minimize computational costs. The frame sampling process is defined as:
$S=\{v_{i}\mid i=k\cdot l+1,\;k\in\mathbb{Z}_{\geq 0},\;k\cdot l\leq m\}$ (1)
where $S$ represents the set of sampled frames, $v_{i}$ denotes the $i$-th frame in the video $V$, $m$ is the total number of frames, and $l$ is the sampling interval. This approach reduces redundancy while ensuring a diverse representation of sampled frames.
From Table 1, it is evident that when the sampling interval $l$ is set to 10, 20, 30, 40, 50, or 60, the EVQAScore remains stable, consistently outperforming the EMScore, PAC-S, and PAC-S++ metrics. This suggests that frame sampling does not significantly influence the EVQAScore calculation. In our experiment, we select a sampling interval of $l=30$, which reduces computational costs without any degradation in performance.
After sampling the frames, we extract fine-grained patch features for more detailed matching. Additionally, we perform keyword extraction to enhance semantic understanding.
Key Object Extraction
For each frame $v_{i}$ in $S$, we use YOLO to extract key objects $O_{i}$, saying:
$O_{i}=\text{YOLO}(v_{i})$ (2)
where $O_{i}=\{o_{i,j}\}_{j=1}^{|O_{i}|}$ stands for a list of object regions in a frame. Each object $o_{i,j}$ is a region in the original video frame $v_{i}$ containing meaningful items. Then we concat all $O_{i}$ to get the final object sequence:
$O=\text{concat}_{i=1}^{|S|}(O_{i})$ (3)
Figure 3: Prompt used for keywords extraction.
@@ 3.2 Data Preparation for EVQAScore
Given a text input $X$, we use a large language model (LLM) to extract a list of key phrases $K=\{k_{1},k_{2},\ldots,k_{n}\}$ as follows:
@@
where $\text{LLM}_{\text{extract}}$ represents the extraction function of the language model, and $P$ denotes the prompt designed to guide the extraction of relevant phrases from the input $X$. This approach improves the understanding of caption semantics and is particularly effective for question answering, especially when the caption is lengthy.
3.3 EVQAScore for Video-Caption Data Quality Evaluation
The improved scoring method now combines the coarse-grained score and the modified fine-grained score, which includes fine-grained feature and keyword-based matching. This enhanced PAC-Score improves efficiency when handling long videos and complex text descriptions, focusing on the most essential frames and phrases.
@@ 3.3 EVQAScore for Video-Caption Data Quality Evaluation
$P(V,X)_{f}=P\left(O,K\right)_{f}=\frac{1}{|X|}\sum_{k_{j}\in K}\max_{o_{i}\in O}f_{o_{i}}^{\top}f_{k_{j}}$
$R(V,X)_{f}=R\left(O,K\right)_{f}=\frac{1}{|O|}\sum_{o_{i}\in O}\max_{k_{j}\in K}f_{o_{i}}^{\top}f_{k_{j}}$
where $f_{o_{i}}$ and $f_{k_{j}}$ represent the CLIP embedding scores of object $o_{i}$ and keyword $k_{j}$. Subsequently, the fine-grained EVQAScore is calculated as:
@@
Finally, the overall EVQAScore is computed by combining the original coarse-grained score with the updated fine-grained score, which incorporates both fine-grained and coarse-grained matching.
@@
This formula ensures that both visual and textual representations are aligned more effectively, improving precision, recall, and computational efficiency by focusing on fine-grained patches and keywords.
3.4 EVQAScore for VideoQA Data Quality Evaluation
We compute the EVQAScore for VideoQA data using the following formula:
$\text{EVQAScore}(\text{Concat}(Q,A),V)$
In this formulation, $\text{Concat}(Q,A)$ represents the concatenation of the question $Q$ and the answer $A$. By employing a keyword extraction technique, our EVQAScore is able to accurately capture the semantics of the QA pair.
Table 1: We compare the time required to evaluate the VATEX-EVAL video-caption dataset. EVQAScore consistently outperforms the baselines, while maintaining an acceptable computational cost. The evaluation is conducted using 8 NVIDIA H20 GPUs.